OpenStudy (anonymous):

rotation about the x-axis : area between curve y=(2+x)^2 , x=-4,x=-3 , i found 31pi/5

ganeshie8 (ganeshie8):

correct. i got 7/3

OpenStudy (anonymous):

thnx

ganeshie8 (ganeshie8):

can you pls tell how you calculated ? i used calculus..

OpenStudy (cwrw238):

this is integral of pi (2 + x)^4 dx between -4 and -3

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

@ganeshie8 the same that @cwrw238 did

ganeshie8 (ganeshie8):

i just did took integral of y from -4 to -3. how pi is coming in between.. can you pls refer me to some doc.. or explain... i fogot my basics.. dont remember much... !!

ganeshie8 (ganeshie8):

@cwrw238

OpenStudy (anonymous):

http://www.twiddla.com/865824 wht cwrw238 is doing he is trying to find out the vo,ume of the rotation

OpenStudy (anonymous):

@ganeshie8 http://www.saylor.org/courses/ma102 Unit 2

OpenStudy (anonymous):

volume

OpenStudy (cwrw238):

=[ pi (2 + x)^5 / 5] = pi (2-4)^5 / 5 - pi (2-3)^5 / 5

OpenStudy (anonymous):

http://www.twiddla.com/865824

ganeshie8 (ganeshie8):

oh yeah... got it. thanks :P

OpenStudy (anonymous):

i did the solution here

ganeshie8 (ganeshie8):

thanks @Nick_Black .. same i did. calculted the area... instead of calculating the volume !! i got ur point thanks :)

OpenStudy (anonymous):

ur welcome ganeshie i am Narayan

OpenStudy (kropot72):

The volume of the solid generated is\[\frac{7 \pi}{3}\]